OpenStudy (anonymous):

Write the equation of the line with a slope of 3/7 passing through the point (9,-2) in standard or slope-point form?

OpenStudy (anonymous):

slope-point would be the easiest. it looks like this: \[(y-y _{1})=m(x-x _{1})\] so all you do is plug in your slope (which is m) and the x and y of your given point for \[x _{1} and y _{1^{?}}\]

OpenStudy (anonymous):

So...: \[(y--2)=3/7(x-9)\] But when I begin to simplify this the numbers become weird... I get \[X_{1} = -27/7\] my answers are: a. 2x-9y=-3/7 b. y=3/7x+9 c. y=3/7x-2 d. 3x-7y=41 Thanks for your help

OpenStudy (anonymous):

None of those answers are in point-slope form... did you write your question correctly? Regardless: Standard form of a line is Ax+By=C. Using your given information, start by writing your line equation in slope intercept (y=mx+b). Rearrange until you have your x+y on one side. Make sure there are no fractions by multiplying to cancel out any denominators. Try it out and let me know if this gets you one of your answers.

OpenStudy (anonymous):

Sorry meant standard and slope-intercept form. When I rearranged the problem to slope-intercept form, I got y=3/7x-8. (Got the -8 by putting in the point (9,-2) to the problem to find my y intercept). Am I doing the problem wrong still..?

OpenStudy (anonymous):

Yeah, -8 is wrong. Try that one more time.

OpenStudy (anonymous):

Ok -8 is wrong... when I correctly put in (9,-2) I get -2=3/7(9)+b and solve for b iget 13/7 which is of course also a weird number... thank you for your patience :)

OpenStudy (anonymous):

mm... check your math one more time! 9*(3/7) is (27/7) right? (-2)-(27/7) is...

OpenStudy (anonymous):

ahah! an elementary mistake. I added the 27/7 instead of subtracted.. still (-2)-(27/7) is -41/7

OpenStudy (anonymous):

yay! okay, now rearrange y=mx+b to make Ax+By=C

OpenStudy (anonymous):

ok so....-3x+y=41? (Multiplied the 7s to get rid of em) so close to answer D...!